Solve the system algebraically. Check your work.
4x + 2y = 10
3x + 2y = 6

Answer by stanbon(75887)

Solve the system algebraically. Check your work.
4x + 2y = 10
3x + 2y = 6
----
Subtract the 2nd equation from the 1st and solve for "x":
x = 4
---
Substitute for "x" and solve for "y":
3(4) + 2y = 6
2y = -6
y = -3
==============
Cheers,
Stan H.
==============